A normal population has a mean μ =15 and standard deviation = 4. What is the probability that a randomly chosen value will be greater than 20?
Answer by stanbon(75887)

A normal population has a mean μ =15 and standard deviation = 4.
What is the probability that a randomly chosen value will be greater than 20?
----
z(20) = (20-15)/4 = 5/4
So, 20 is (5/4)th standard deviations above the mean.
------
P(x > 20) = P(z > 5/4) = normalcdf(5/4,100) = 0.1056
------
Cheers,
Stan H.
----------